OpenStudy (anonymous):

Medal to help _The area of a rectangular pool is 10x 2 - 19x-15 the length of the poll is 5x +3..What is the width ? a. 2x -18 b.2x-5 c.5x-5 d.5x-22 I think it is b.2x-5

OpenStudy (raffle_snaffle):

well multiply the length by each choice

OpenStudy (raffle_snaffle):

you are right

OpenStudy (anonymous):

The area is length*width. This allows us to check your answer: \[ \text{length * your width = }(5x + 3)(2x - 5) = 10x^2 - 25x + 6x - 15\\ = 10x^2 - 19x - 15 \text{ = area}\]